# NeuroBench: A Framework for Benchmarking Neuromorphic Computing Algorithms and Systems

### v3

@@ Abstract
Neuromorphic computing shows promise for advancing computing efficiency and capabilities of AI applications using brain-inspired principles. However, the neuromorphic research field currently lacks standardized benchmarks, making it difficult to accurately measure technological advancements, compare performance with conventional methods, and identify promising future research directions. Prior neuromorphic computing benchmark efforts have not seen widespread adoption due to a lack of inclusive, actionable, and iterative benchmark design and guidelines. To address these shortcomings, we present NeuroBench: a benchmark framework for neuromorphic computing algorithms and systems. NeuroBench is a collaboratively-designed effort from an open community of nearly 100 co-authors across over 50 institutions in industry and academia, aiming to provide a representative structure for standardizing the evaluation of neuromorphic approaches. The NeuroBench framework introduces a common set of tools and systematic methodology for inclusive benchmark measurement, delivering an objective reference framework for quantifying neuromorphic approaches in both hardware-independent (algorithm track) and hardware-dependent (system track) settings. In this article, we present initial performance baselines across various model architectures on the algorithm track and outline the system track benchmark tasks and guidelines. NeuroBench is intended to continually expand its benchmarks and features to foster and track the progress made by the research community.
@@ Introduction
As Figure 1 shows, the NeuroBench framework involves two tracks to enable agile algorithm and system development. As an emerging technology, neuromorphic hardware has not converged to a single platform which is commercially available, thus a large fraction of neuromorphic research explores algorithmic advancement on conventional systems which may not be optimal for performance. Thus, NeuroBench consists of an algorithm track for hardware-independent evaluation and a system track for fully deployed solutions. The algorithm track defines four novel benchmarks for neuromorphic methods across diverse domains, namely few-shot continual learning, computer vision, motor cortical decoding, and chaotic forecasting, and utilizes complexity metrics to analyze solution costs. Such hardware-independent benchmarking enables algorithmic exploration and prototyping, especially when simulating algorithm execution on non-neuromorphic platforms. Meanwhile, the system track defines standard protocols to measure the real-world speed and efficiency of neuromorphic hardware on benchmarks ranging from standard machine learning tasks to promising fields for neuromorphic systems, such as optimization. Both the algorithm and system track will be extended and co-developed as NeuroBench continues to expand.
The following Results section organizes descriptions of the algorithm track benchmark framework and its baseline results, as well as specifications of the system track benchmark framework and tasks. Further details regarding the benchmark metric formulations, task specifications, and baseline solutions can be found in the Methods section. Baseline results on NeuroBench benchmarks outline unexplored research opportunities in optimizing algorithmic architectures and training of sparse, stateful models to achieve greater performance and resource efficiency. As NeuroBench is intended to continually grow over time, the latest developments and opportunities to engage with the project are reported on the website.aaahttps://neurobench.ai
Results
@@
The complete NeuroBench framework is shown in Figure 1. It includes two tracks with defined datasets, metrics, and modular evaluation components to enable flexible development. The algorithm track focuses on hardware-independent algorithm prototyping to identify promising methods. These in turn inform system design by highlighting target algorithms for optimization and relevant system workloads for benchmarking. The system track enables optimization and evaluation of performant implementations, providing feedback to refine algorithmic complexity modeling and analysis. The interplay between the tracks creates a virtuous cycle: algorithm innovations guide system implementation, while system-level insights accelerate further algorithmic progress. This approach allows NeuroBench to advance neuromorphic algorithm-system co-design.
@@
In the next few sections, we describe the algorithm track, including general complexity metric definitions, benchmark tasks, and common infrastructure tooling. We apply the framework to report baseline results for each algorithm benchmark. Then, we specify protocols and tasks established in the system track to assess deployed neuromorphic performance across promising application workloads. By outlining both tracks, we provide a roadmap towards standardizing benchmark procedures in both hardware-independent and hardware-dependent settings.
@@
  • Model Execution Rate – Execution rate, in Hz, of the model computation based on forward inference passes per second, measured in the time-stepped simulation timescale. The time is correlated to real-world data time. For example, if a model is designed to process data from an event camera with 50 ms input stride, the model execution rate is 20 Hz. This metric provides intuition into the deployed real-time responsiveness of a model, as well as its computational requirements.
@@
Footprint and connection sparsity are classified as static metrics, which can be analytically determined from the model only. Activation sparsity, synaptic operations, and correctness are classified as workload metrics, which are dependent on execution or simulation of the model based on the benchmark data. Model execution rate is an exception, as it is a feature of the algorithm which neither needs to be calculated nor extracted from the model or its outputs, and thus is reported directly by the solution designer in benchmark results.
@@
Task Dataset Correctness metric
Task description
Keyword FSCIL MSWC [21] Accuracy
Few-shot, continual learning of keyword classes.
Event Camera Object Detection Prophesee 1MP Automotive [22] COCO mAP
Detecting automotive objects from event camera video.
[…]
Predicting fingertip velocity from cortical recordings.
Chaotic Function Prediction Mackey-Glass time series [24] sMAPE
Autoregressive modeling of chaotic functions.
@@
The NeuroBench algorithm benchmarks are wrapped in a harness which standardizes the benchmark interfaces. The harness provides benchmark users with a consistent framework for loading data, processing data and model outputs, and calculating and reporting metrics, thereby ensuring fair and standard comparisons of the results. It is built with straightforward interfaces which are designed to be extended with new frameworks, algorithms, and tasks. The benchmark harness is open-source for use and development.bbbhttps://github.com/NeuroBench/neurobench
@@
Currently, the harness and all baseline models are built using PyTorch [34] or frameworks based on it, such as snnTorch [18] and SpikingJelly [35]. Due to its modular structure and simple interfaces, the harness can grow to be compatible with further neuromorphic tools such as Lava [36] and Fugu [37]. Furthermore, it also supports the extension of data and metric pipelines in order to implement additional benchmark tasks. Benchmarks outside of the NeuroBench v1.0 suite can make use of the harness infrastructure for open reproducibility, and also to garner interest in the community towards inclusion in a future NeuroBench version, through which the task will have long-term support and appear in affiliated leaderboards and workshop events.
@@
Furthermore, the synaptic operations metric, intended to capture model computation cost, currently does not account for neuron updates. The dynamics of neuron models, including mechanisms like leakage and reset, can vary heavily in complexity. However, counting the number and type of operations from neuron updates, as well as estimating their overall costs, depends on the specific arithmetic or circuit implementation. Thus, they are not accounted for in the broader algorithmic complexity metrics. Solution-specific metrics that assume a particular implementation platform, as have been defined previously [38], can be used to estimate neuron update costs. These estimates can then be combined with the total number of neuron updates per model computation to measure overall neuron operation complexity during evaluation.
@@
[…]
@@
While the algorithm track aims to benchmark solutions in a system-independent manner via complexity analysis, the NeuroBench system track aims to evaluate deployed latency, throughput, and efficiency of systems comprised of an algorithm deployed to a hardware platform. In order for the hallmarks of neuromorphic hardware to be aptly judged against conventional systems and foster the expansion of neuromorphic solutions, fair comparisons must be made between sufficiently mature neuromorphic systems and conventional systems solving the same tasks.
@@
A key challenge for benchmarking neuromorphic hardware is that systems are implemented and deployed at vastly different scales to serve diverse applications, from cloud services (e.g. multi-chip platforms like Loihi [50] and SpiNNaker [51]) to embedded sensing intelligence (e.g. Speck [52] and SNP [53, 54]). This range is visualized in Figure 6. Existing benchmarks for conventional machine learning systems separate submissions between datacenter-level computing [19] and embedded processing [55]. Thus, rather than pursuing a one-size-fits-all suite of tasks for neuromorphic systems, the goal of the NeuroBench system track is to develop benchmarks at various scales and use cases, united under a common set of guidelines and measurement methodology, thereby embracing the diversity of neuromorphic approaches.
In this section, we present the system track guidelines outlining the metrics, tasks, and harness components, representing collective design between multiple owners and vendors of neuromorphic hardware. As the system track benchmark tooling is currently under development, insight from the algorithm track’s early results will be exploited towards a future release of the detailed harness documentation, benchmark procedure, and baselines for the system track by the end of 2024.
@@
In order to be representative of the properties of a deployed system, the system benchmarks, like the algorithm benchmarks, are assessed at the task level for the overall system, as opposed to operation or kernel level assessment of individual components. Task-level benchmarks enable straightforward comparison between systems of any type with regard to their abilities to solve problems, and the overall system-level measurement describes the realistic capability and efficiency of a whole solution. For ease of comparison and benchmark result analysis, key features of the NeuroBench system track are consistent with features from the widely-adopted machine learning system benchmark MLPerf Inference [19].
Task Scenarios.
Where applicable, the NeuroBench system track will utilize benchmark task scenarios from MLPerf. MLPerf describes task scenarios under which system benchmarks are presented to the system under test (SUT). For large-scale batched processing, MLPerf defines the Offline and Server scenarios, which are latency-unconstrained and latency-constrained, respectively, with performance measured in terms of prediction throughput. For batch-size-1 processing, MLPerf defines Single-stream scenarios, for which performance is measured in terms of prediction latency.
However, the MLPerf task scenarios all consider data as discrete samples, and therefore do not fit with some key neuromorphic system applications. The NeuroBench system track thus defines two additional task scenarios: Real-time and Optimization. Real-time uses continuous data streams (e.g., from an event camera), which must be processed online by the SUT, in contrast with the Single-stream scenario, where samples are sent only once the SUT has completed processing. Optimization applications use heuristic methods for otherwise intractable problems, and thus do not have notions of sample throughput or sample latency. The Optimization scenario benchmarks will report latency for the SUT to reach multiple correctness thresholds, as many approaches find sucessively improved solutions over time.
For each task and scenario, the following metric categories are reported:
@@
  • Correctness – Due to the tight coupling between an algorithm and its system implementation in many existing neuromorphic hardware solutions, the particular model used to solve the benchmark task is unconstrained. Therefore, correctness must be measured to verify the validity of the solution. No correctness thresholds are imposed on submissions, but the benchmark leaderboard will impose tiers of solution correctness on submissions to evaluate accuracy-efficiency trade-offs of system approaches.
  • Performance – Depending on the task scenario, the performance of the system is measured differently. The Offline and Server scenarios report throughput, while the Single-stream scenario reports latency. The Real-time scenario similarly reports latency, which for a large portion of execution (e.g., 90%, defined per-benchmark) must not exceed a real-time threshold, or else the submission is not considered valid. The Optimization scenario reports the time to solution at various solution quality thresholds, defined per-benchmark.
@@
  • Efficiency – Conventional system benchmarks such as TOP500 [56] for HPC and MLPerf Inference [19] for deep learning do not require power measurement submission in the main benchmark, instead allowing for separate submissions to an adjacent power track (Green500 [57] and MLPerf Power [58], respectively). Not only has efficiency been usually considered as a second-order metric for conventional systems, it is also notoriously difficult to precisely measure. However, as energy efficiency is a key hallmark of biology and thus is a focus of neuromorphic research, power and energy consumption must be first-order metrics in the NeuroBench system track.
    The Server, Offline, and Real-time scenarios report average power as their batched and online processing is continuous, the Single-stream scenario reports the energy per-sample, and the Optimization scenario reports total energy consumed at each solution quality threshold.
@@
Benchmark submissions may perform separate runs to report performance and power in order to demonstrate system flexibility (e.g., a ‘performance-mode’ run optimal for latency and an ‘efficiency-mode’ run optimal for energy), however in all runs, both metrics must be reported.
@@
Importantly for the NeuroBench system track, in measuring latency and efficiency, data pre- and post-processing must be taken into account. Neuromorphic methods will often consume and produce non-standard (e.g., event-based) data modalities, the processing of which may consume a significant amount of the overall execution latency and may not be computed on the neuromorphic hardware itself. As many instances of neuromorphic hardware cannot be deployed without such associated processing, it is essential that latency timing and efficiency measurement captures the cost of data processing, which stands in contrast with conventional system benchmarks that measure starting from pre-processed data [59].
@@
Benchmarks for the system track will include tasks of interest to neuromorphic systems, from embedded to datacenter scales. The tasks are key application areas for existing systems, and they differ from the tasks in the algorithm track, which are more research-oriented. Towards future iterations of the NeuroBench suite, the algorithm and system tracks are intended to coalesce as both algorithms and systems mature. As the benchmark results identify properties of highly effective algorithms and systems, the algorithm and system tracks will converge to the same selection of tasks that are seen as the most impactful for future progress in the field. Two benchmark specifications for the system track are defined in this article.
@@
  • Acoustic Scene Classification – The acoustic scene classification benchmark challenges systems to classify audio into predefined categories based on the environmental audio context. Such capabilities are key for hearable devices, which can utilize them to automatically adjust sound equalisation profiles, appropriately target microphone denoising, and support active noise cancellation. The application further challenges systems to fulfill technical requirements, such as always-on and real-time operation, and time series processing. Acoustic scenes provide a rich repertoire of features that are necessary for prediction, thus this tasks is a complement to keyword classification, which mainly focuses on shorter-term features (e.g., phonemes) with a relatively smaller feature repertoire.
@@
    The benchmark evaluates the classification capabilities of both neuromorphic systems and conventional computing platforms using datasets such as those from the DCASE challenge [60] (if permissible, subject to license). These datasets consist of a myriad of audio recordings from diverse environments, including airports, public parks, and buses, thus providing a comprehensive foundation for testing both application- and system-level performance.
    The task will be presented under the Real-time or Single-stream task scenarios, providing a continuous audio stream or sliced samples to the SUT, during which the acoustic scene periodically changes. Classification probability will be sampled to determine the correctness of the prediction. The average power measured during inference is a key indicator of the efficiency and always-on capability of the SUT, and inference latency will be measured in terms of the prediction time relative to the onset of each new acoustic scene.
@@
  • QUBO – As an Optimization scenario task, NeuroBench incorporates quadratic unconstrained binary optimization (QUBO). QUBO is a particularly beneficial first optimization task for NeuroBench for two reasons. First, the binary variables are inclusive to neuromorphic systems with purely binary spike communication. Second, real-world QUBO applications typically feature sparse cost matrices [61] which benefit from the sparse synaptic connectivity / matrix multiplication that neuromorphic systems are often optimized for [62]. The initial set of QUBO workloads in NeuroBench searches for the maximum independent (i.e. unconnected) set of nodes in graphs, a task that has wide applications across industry and academia, such as resource allocation in wireless networks, portfolio optimization, and task scheduling.

### v5

@@ Abstract
Neuromorphic computing shows promise for advancing computing efficiency and capabilities of AI applications using brain-inspired principles. However, the neuromorphic research field currently lacks standardized benchmarks, making it difficult to accurately measure technological advancements, compare performance with conventional methods, and identify promising future research directions. Prior neuromorphic computing benchmark efforts have not seen widespread adoption due to a lack of inclusive, actionable, and iterative benchmark design and guidelines. To address these shortcomings, we present NeuroBench: a benchmark framework for neuromorphic computing algorithms and systems. NeuroBench is a collaboratively-designed effort from an open community of researchers across industry and academia, aiming to provide a representative structure for standardizing the evaluation of neuromorphic approaches. The NeuroBench framework introduces a common set of tools and systematic methodology for inclusive benchmark measurement, delivering an objective reference framework for quantifying neuromorphic approaches in both hardware-independent (algorithm track) and hardware-dependent (system track) settings. In this article, we outline tasks and guidelines for benchmarks across multiple application domains, and present initial performance baselines across neuromorphic and conventional approaches for both benchmark tracks. NeuroBench is intended to continually expand its benchmarks and features to foster and track the progress made by the research community.
@@ Introduction
As Figure 1 shows, the NeuroBench framework involves two tracks to enable agile algorithm and system development. As an emerging technology, neuromorphic hardware has not converged to a single platform which is commercially available, thus a large fraction of neuromorphic research explores algorithmic advancement on conventional systems which may not be optimal for performance. Thus, NeuroBench consists of an algorithm track for hardware-independent evaluation and a system track for fully deployed solutions. The algorithm track defines four novel benchmarks for neuromorphic methods across diverse domains, namely few-shot continual learning, computer vision, motor cortical decoding, and chaotic forecasting, and utilizes complexity metrics to analyze solution costs. Such hardware-independent benchmarking enables algorithmic exploration and prototyping, especially when simulating algorithm execution on non-neuromorphic platforms. Meanwhile, the system track defines standard protocols to measure the real-world speed and efficiency of neuromorphic hardware on benchmarks ranging from standard machine learning tasks to promising fields for neuromorphic systems, such as optimization.
@@
Each NeuroBench track includes defined datasets, metric and measurement methodology, and modular evaluation components to enable flexible development. Promising methods identified from the algorithm track will inform system design by highlighting target algorithms for optimization and relevant system workloads for benchmarking. The system track in turn enables optimization and evaluation of performant implementations, providing feedback to refine algorithmic complexity modeling and analysis. The interplay between the tracks creates a virtuous cycle: algorithm innovations guide system implementation, while system-level insights accelerate further algorithmic progress. This approach allows NeuroBench to advance neuromorphic algorithm-system co-design. Both the algorithm and system track will be extended and co-developed as NeuroBench continues to expand.
@@
In the next few sections, we describe the algorithm track, including general complexity metric definitions, benchmark tasks, and common infrastructure tooling. We apply the framework to report baseline results for each algorithm benchmark, which outline unexplored research opportunities in optimizing algorithmic architectures and training of sparse, stateful models to achieve greater performance and resource efficiency. Then, we show baseline results established in the system track to assess neuromorphic performance across promising application workloads. By outlining both tracks, we provide a roadmap towards standardizing benchmark procedures in both hardware-independent and hardware dependent settings.
@@
Footprint and connection sparsity are classified as static metrics, which can be analytically determined from the model only. Activation sparsity, synaptic operations, and correctness are classified as workload metrics, which are dependent on execution or simulation of the model based on the benchmark data.
In addition to the above complexity metrics, the algorithm track proposes to define Model Execution Rate, corresponding to the rate, in Hz, at which the model’s forward inference pass needs to be executed. For example, if a model is designed to process data from an event camera with a 50 ms input stride, the model execution rate is 20 Hz. The execution rate is a critical feature of the algorithm which provides intuition into the tradeoff between latency and computational footprint of a deployed model, and is reported directly by the solution designer in benchmark results since it neither needs to be calculated nor extracted from the model or its outputs.
@@
Task Dataset Correctness metric Task description
Keyword FSCIL MSWC [21] Accuracy Few-shot, continual learning of keyword classes.
Event Camera Object Detection Prophesee 1MP Automotive [22] COCO mAP Detecting automotive objects from event camera video.
NHP Motor Prediction Primate Reaching [23] R2 Predicting fingertip velocity from cortical recordings.
Chaotic Function Prediction Mackey-Glass time series [24] sMAPE Autoregressive modeling of chaotic functions.
@@
The NeuroBench algorithm benchmarks are wrapped in a harness which standardizes the benchmark interfaces. The harness provides benchmark users with a consistent framework for loading data, processing data and model outputs, and calculating and reporting metrics, thereby ensuring fair and standard comparisons of the results. It is built with straightforward interfaces which are designed to be extended with new frameworks, algorithms, and tasks. The benchmark harness is open-source for use and development (https://github.com/NeuroBench/neurobench).
@@
Currently, the harness and all baseline models are built using PyTorch [34] or frameworks based on it, such as snnTorch [18] and SpikingJelly [35]. Due to its modular structure and simple interfaces, the harness can grow to be compatible with further neuromorphic tools such as Lava [36] and Fugu [37]. Furthermore, it also supports the extension of data and metric pipelines in order to implement additional benchmark tasks. Widely validated benchmarks in keyword [38] and gesture classification [12], which are foundational in neuromorphic and conventional research [39, 13, 40, 41], have been incorporated into the harness to complement the novel tasks in the NeuroBench v1.0 suite. Any novel or existing benchmarks can make use of the harness infrastructure for open reproducibility, and also to garner interest in the community towards long-term task support and appearance in NeuroBench-affiliated leaderboards and challenge events.
@@
Furthermore, the synaptic operations metric, intended to capture model computation cost, currently does not account for neuron updates. The dynamics of neuron models, including mechanisms like leakage and reset, can vary heavily in complexity. However, counting the number and type of operations from neuron updates, as well as estimating their overall costs, depends on the specific arithmetic or circuit implementation. Thus, they are not accounted for in the broader algorithmic complexity metrics. The algorithmic metric framework can be extended with solution-specific metrics that assume a particular implementation platform to estimate neuron update costs, which have been previously defined [42]. These estimates can then be combined with the total number of neuron updates per model computation to measure overall network operation complexity during evaluation.
@@
Figure 3: Test accuracy per session on the keyword FSCIL task for prototypical and frozen baselines, with the accuracy on both base classes and incrementally-learned classes (left), and accuracy on all incrementally-learned classes only (right). Incremental session 0 refers to the accuracy on base classes after pre-training only. Shaded area represents 5th and 95th percentile on 100 runs. Frozen baselines with no adaptation do not learn incremental classes and thus have a fixed 0% accuracy for New Classes Performance.
@@
While the algorithm track aims to benchmark solutions in a system-independent manner via complexity analysis, the NeuroBench system track aims to evaluate deployed execution time, throughput, and efficiency of systems comprised of an algorithm deployed and tailored to a hardware platform. Previous benchmark studies have examined neuromorphic systems under various applications, including keyword spotting [39, 54], audio and video processing [55], and combinatorial optimization [56, 57]. While these studies have demonstrated neuromorphic system advantages, the benchmark tasks have been unaligned. In order for the hallmarks of neuromorphic hardware to be aptly judged against conventional systems and foster the expansion of neuromorphic solutions, transparent and objective comparisons must be made on standard tasks between sufficiently mature neuromorphic systems head-to-head, as well as against conventional systems.
@@
A key challenge for benchmarking neuromorphic hardware is that systems are implemented and deployed at vastly different scales to serve diverse applications, from cloud services (e.g., multi-chip platforms like Loihi [58] and SpiNNaker [59]) to embedded sensing intelligence (e.g., Speck [60] and SNP [61, 62]). This range is visualized in Figure 6. Existing benchmarks for conventional systems have individual focuses across high-performance [63], datacenter-level computing [19], and embedded processing [64]. Thus, rather than pursuing a one-size-fits-all suite of tasks, the goal of the NeuroBench system track is to develop benchmarks at various scales and use cases, under multiple application areas in which both conventional and neuromorphic platforms may compete. The selected v1.0 NeuroBench system track benchmarks represent key commercial application areas for existing systems, and they differ from the tasks in the algorithm track, which are more research-oriented. As benchmark results continue to identify properties of highly effective algorithms and systems, the two tracks will converge to the same selction of tasks that are seen as the most impactful for future progress in the field.
In this section, we present the system track guidelines outlining metrics and tasks, representing collective design between multiple owners and vendors of neuromorphic hardware. Baseline benchmark results for neuromorphic and conventional systems are reported, and further official results will be collected and announced at a regular cadence, akin to the MLPerf suite [65]. As with all other facets of the NeuroBench framework, the system track guidelines will continue to be adapted and extended iteratively as benchmark results are produced and shared. Up-to-date information on the latest benchmarks and official results can be found on the NeuroBench website (https://neurobench.ai/).
@@
In order to be representative of the properties of a deployed system, the system benchmarks, like the algorithm benchmarks, are assessed at the task level for the overall system, as opposed to operation- or kernel-level assessment of individual components. Task-level benchmarks enable straightforward comparison between systems of any type with regard to their abilities to solve problems, and the overall system-level measurement describes the realistic capability and efficiency of a whole solution.
Each individual system benchmark uses task-specific metrics aligned with correctness, timing, and efficiency to measure the system under test (SUT). The following general considerations are applied to each category:
  • Correctness – In other system benchmarks, such as the closed category of the MLPerf Inference framework [19], the same trained model is used to benchmark all SUTs, and a correctness threshold is imposed to ensure optimizations such as lower precision do not disrupt task performance.
@@
    Due to the tight coupling between an algorithm and its system implementation in many existing neuromorphic hardware solutions, the particular model used to solve a NeuroBench system track benchmark task is unconstrained. Therefore, correctness must be measured to verify the validity of the solution. No correctness thresholds are imposed on submissions, but the benchmark leaderboard will impose tiers of solution correctness on submissions to evaluate accuracy-efficiency trade-offs of system approaches.
  • Timing – Depending on the task, timing performance can include measurements of sample throughput or execution time. Individually, the former entails an offline, batched inference benchmark, while the latter aligns with a streaming benchmark, in which one inference does not start until the previous one ends. Together, both throughput and execution time should be reported for tasks in which the SUT runs multiple inferences at any given time, each representing a request which must be responded to within a constrained window. The MLPerf Inference framework has defined widely-adopted general task scenarios corresponding to each of these categories (offline, single-stream, and server, respectively), and the NeuroBench system track will use these scenario guidelines where applicable to maintain consistency and build on conventional frameworks.
    In addition, neuromorphic systems are also applied to tasks in which there is no notion of discrete sample throughput or execution, such as for heuristic approximations of intractable problems or operation over a continuous stream of data (e.g., from an event camera). Timing performance should be defined on a per-benchmark basis for such tasks, such as a time-to-solution latency or percentage of execution which exceeds a real-time threshold.
@@
  • Efficiency – Conventional system benchmarks such as TOP500 [63] for HPC and MLPerf Inference [19] for deep learning do not require power measurement submission in the main benchmark, instead allowing for separate submissions to an adjacent power track (Green500 [66] and MLPerf Power [67], respectively). Not only has efficiency been usually considered as a second-order metric for conventional systems, it is also notoriously difficult to precisely measure. However, as energy efficiency is a key hallmark of biology and thus is a focus of neuromorphic research, power and energy consumption must be first-order metrics in the NeuroBench system track. Similarly to timing metrics, efficiency metrics should be tailored on a per-benchmark basis, i.e., a real-time always-on processing task may focus on average power, while offline batched systems focusing on high-throughput inference may focus on both peak power and energy per inference.
As neuromorphic systems currently explore a broad range of varied implementation approaches, board-level integration, and developmental stages of hardware, platform diversity poses difficult challenges for completely consistent hardware measurement methodologies (e.g., consistent power meters, chip interfaces, and data loading). Thus, to enable an initial step towards consistency in the system track while ensuring openness, we focus on the development of guidelines for transparent documentation, as they provide the foundation for shared methodology among highly diverse solutions. While there may be differences in how metrics are measured, salient details will be available to contextualize the results, allowing for holistic analysis, and leading the way for future consistency by enforcing transparency.
@@
Benchmark submissions may perform separate runs to report performance and power in order to demonstrate system flexibility (e.g., a ‘performance-mode’ run optimal for execution time and an ‘efficiency-mode’ run optimal for energy), however in all runs, both metrics must be reported.
@@
Importantly for the NeuroBench system track, in measuring timing and efficiency, data pre- and post-processing must be taken into account. Neuromorphic methods will often consume and produce non-standard (e.g., event-based) data modalities, the processing of which may consume a significant amount of the overall execution time and may not be computed on the neuromorphic hardware itself. As many instances of neuromorphic hardware cannot be deployed without such associated processing, it is essential that measurements capture the cost of data processing, which stands in contrast with conventional system benchmarks whose measurements start from pre-processed data [65].
@@
Two benchmark specifications for the v1.0 system track are defined in this article, covering embedded to datacenter scales. Full benchmark details are available in the Methods section.
@@
  • Acoustic Scene Classification – The acoustic scene classification benchmark challenges systems to classify audio into predefined categories based on the environmental audio context. Such capabilities are key for hearable devices, which can utilize them to automatically adjust sound equalisation profiles, appropriately target microphone denoising, and support active noise cancellation. The application further challenges systems to fulfill technical requirements, such as always-on and real-time operation, and time series processing. Acoustic scenes provide a rich repertoire of features that are necessary for prediction, thus this task is a complement to keyword classification, which mainly focuses on shorter-term features (e.g., phonemes) with a relatively smaller feature repertoire.
@@
    The benchmark evaluates the classification capabilities of both neuromorphic systems and conventional computing platforms using datasets from the DCASE challenge [68]. These datasets consist of a myriad of audio recordings from diverse environments, including airports, public parks, and buses, thus providing a comprehensive foundation for testing both application- and system-level performance. The NeuroBench subset of the DCASE dataset includes 41360/16240 train/test samples across four classes (airport, street traffic, bus, park).
    The task will be presented under the single-stream task scenario, providing one 1-second sample to the SUT at a time. Classification probability will be sampled to determine the correctness of the prediction. As the NeuroBench system track allows for unconstrained algorithmic implementation, pre-processing and inference metrics should be separately measured and reported together, which differs from prior system benchmarking that only measure inference [39, 64]. Timing results report on-device average execution time per sample. Since the platform diversity of edge-targeted systems poses inherent inconsistencies in efficiency measurement, power should be reported under idle and active contexts, following prior benchmark study methodology [39]. Idle power measures the system prepared for inference with the model loaded, and active power measures the system running pre-processing or inference. The difference between active and idle measurements offers dynamic power, which is used along with execution time to calculate dynamic energy-per-sample.
@@
  • QUBO – As a non-ML task, NeuroBench incorporates quadratic unconstrained binary optimization (QUBO). QUBO is a particularly beneficial first optimization task for NeuroBench for two reasons. First, the binary variables are a natural fit for neuromorphic systems with purely binary spike communication. Second, real-world QUBO applications typically feature sparse cost matrices [69] which benefit from the sparse synaptic connectivity and execution that neuromorphic systems are often optimized for [70]. The initial set of QUBO workloads in NeuroBench searches for the maximum independent (i.e. unconnected) set of nodes in graphs, a task that has wide applications across industry and academia, such as resource allocation in wireless networks, portfolio optimization, and task scheduling [71].